OpenStudy (anonymous):

I am not understanding these types of problems. 27-a^3b^6 Help!

OpenStudy (anonymous):

what you have to do there? ...

OpenStudy (anonymous):

maybe.. \[\LARGE 27-a^3b^6=3^3-(ab^2)^3=...\]

OpenStudy (campbell_st):

it looks like the difference of cubics... \[27 = 3^3\] \[a^3b^6 = (ab^2)^3\] \[(3-ab^2)(9 + 3ab^2 + a^2b^4)\]

OpenStudy (anonymous):

yeah.. it looks like. :)

OpenStudy (anonymous):

Yeah I thought it was difference of cubes but I can't seem to factor it correctly once I use the formula.

OpenStudy (anonymous):

Can someone walk me through it. I have a ton of these problems. And I am doing something wrong.

OpenStudy (anonymous):

post new question on wall, not here :)